Militza1720:

A homeowner wants to fence in a rectangular yard along one side of her house. The total area she can fence is 800 sq. ft 2 sides are x and one side is 60-x What are the possible values for x that results in using the least amount of fencing around the space And what is the least amount of fencing needed to create the 800 square foot area

Aricuzyas:

@AZ

Spectrum:

@AZ

carmelle:

@AZ

Militza1720:

This is for algebra class

AZ:

So you're not taking calculus?

Militza1720:

No we’re working on quadratic equations

AZ:

I see, okay so this is our diagram 

Militza1720:

1 attachment
AZ:

we only have three sides to fence so the total material we need is x + x + 60-x and we know that the total area should be 800 ft^2

AZ:

do you know what the formula is for the area of a circle?

Militza1720:

Total area to fence 800 square feet

AZ:

we have a rectangle here with one side as 'x' and the other side as (60-x) what would the area be?

Militza1720:

x(60-x)

AZ:

Good! And remember we said that the area is going to be 800 feet so x(60-x) = 800 can you multiply the x inside?

AZ:

distribute the x into (60-x)

Militza1720:

60x-x^2?

AZ:

Good! so we have 60x - x^2 = 800 what do you get if you subtract 800 on both sides?

Militza1720:

60x-x^2-800=0

AZ:

Good and we can rearrange that so we have -x^2 +60x - 800 = 0 and we can solve it using the quadratic formula \( x = \dfrac{ -b \pm \sqrt{b^2 - 4ac}}{2a} \) when the equation is \( ax^2 + bx + c = 0\)

AZ:

were you able to plug it into the quadratic formula? @Militza1720

Militza1720:

Yes did that and I got x=20 or 40

AZ:

Good! So now let's see which value gives us the smallest perimeter because then that means we've used the least amount of material for fencing so remember perimeter = x + x + x - 60 perimeter = 3x - 60 do you get a smaller perimeter when x is 20 or when x is 40?

Militza1720:

20

AZ:

what is the perimeter when x is 20? that's the answer to your second question

Militza1720:

Would it equal 0?

AZ:

No, think about it. Would it make sense to require 0 feet of material for a fence with an area of 800? Fencing material needed = perimeter = 3x - 60 plug in 20 3*(20) + 60 =

Militza1720:

120

Militza1720:

So why does -60 turn into +60

AZ:

oh I see what you mean

AZ:

my mistake, I accidentally wrote + instead of - so yeah when we plug in x = 20 we see that it doesn't work out so that means the only solution is x = 40 so plug in 40

Militza1720:

Ok so I got 60

AZ:

wait I was doing the perimeter all wrong, I'm sorry

AZ:

perimeter = x + x + (60 - x) perimeter = total fencing material = x + x + 60 - x = x +60

AZ:

I was just going off of memory so I accidentally switched the signs but do you understand now? so x=20 would work and that would give you the smallest perimeter

Militza1720:

Awesome thank you now I got it

AZ:

No problem! Sorry for my careless errors :)

Militza1720:

No thank you for being thorough

AZ:

Of course, it was my pleasure :)